What's the proportion of x+3/4 = 7/8
Found 2 solutions by rfer, youngmike150:
Answer by rfer(16322)

x=7/8-3/4
x=7/8-6/8
x=1/8

Answer by youngmike150(1)

11 and 12 is equivalent to ?